OpenStudy (anonymous):

A light wave travels at a speed of 3.0 × 10^8 meters/second. If the wavelength is 7.0 × 10^-7 meters, what is the frequency of the wave?

OpenStudy (anonymous):

so wavelength = V/F. So you go 7.0*10^-7=(3.0*10^8)/frequency, then solve for frequency. I get 4.29*10^14

OpenStudy (anonymous):

but you should double check that

OpenStudy (anonymous):

yea it was correct:) thanks

OpenStudy (anonymous):

Sure thing!